OpenStudy (anonymous):

What should be added to both sides of this equation to solve for the variable?18 + m = 27

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

so is it just ganna be -18

OpenStudy (anonymous):

18+m=27 -18 -18 m= 19

OpenStudy (anonymous):

i send u another one by message